# Optimal Partial Graph Matching

### v2

@@ Abstract
Partial graph matching addresses the limitations of traditional graph matching by allowing some nodes to remain unmatched, making it applicable to more complex scenarios. However, this flexibility introduces additional complexity, as both the subset of nodes to match and the optimal mapping must be determined. While recent studies have explored deep learning techniques for partial graph matching, a significant limitation remains: the absence of an optimization objective that fully captures the problem’s intrinsic nature while enabling efficient solutions. In this paper, we propose a novel optimization framework for partial graph matching, inspired by optimal partial transport. Our approach formulates an objective that enables partial assignments while incorporating matching biases, using weighted total variation as the divergence function to guarantee optimal partial assignments. We employ the Hungarian algorithm to achieve efficient, exact solutions with cubic time complexity. Our contributions are threefold: (i) we introduce a robust optimization objective that balances matched and unmatched nodes; (ii) we establish a connection between partial graph matching and the linear sum assignment problem, enabling efficient solutions; (iii) we propose a deep graph matching architecture with a novel partial matching loss, providing an end-to-end solution. The empirical evaluations on standard graph matching benchmarks demonstrate the efficacy of the proposed approach.
@@ 1 Introduction
Graph matching is a fundamental problem in network analysis, aiming to establish one-to-one correspondences between nodes in two graphs based on a defined objective. It has broad applications across various fields, including computer vision (Sun et al., 2020), bioinformatics (Zaslavskiy et al., 2009), and social network analysis (Zhang et al., 2019), where it is used to solve complex real-world problems. Traditional graph matching assumes a bijective mapping between the nodes of two graphs or a total injective mapping from the smaller graph to the larger one. However, these assumptions often restrict its applicability in more complex, real-world scenarios. Partial graph matching, a generalized version of the graph matching problem, addresses these limitations by allowing some nodes in both graphs to remain unmatched (Wang et al., 2023a; Jiang et al., 2022). It seeks an optimal partial assignment, an injective partial function between node sets, thereby expanding the practical utility of graph matching. This approach is particularly useful in applications where not all nodes have meaningful counterparts. For example, in image keypoint matching, not all keypoints in two given images have correspondences (Jiang et al., 2022), and in biological networks, certain proteins may lack direct counterparts in other species (Zaslavskiy, 2010).
@@
Solving partial graph matching is inherently more complex than traditional graph matching due to the additional challenge of determining both the subset of nodes to match and the optimal mapping itself. Graph matching is generally framed as a combinatorial optimization problem, where the choice of optimization objective is crucial. Early works formulated it as a Quadratic Assignment Problem (QAP), which is NP-hard (Koopmans & Beckmann, 1957; Lawler, 1963). Recent approaches (Wang et al., 2021; 2020a; Rolínek et al., 2020) use neural networks to learn node representations and derive a cross-graph node-to-node cost matrix. This allows graph matching to be framed as a linear sum assignment problem (Burkard et al., 2012; Yu et al., 2020), assuming a bijective mapping between two node sets of equal size. Even if two graphs have an unequal number of nodes, the problem can still be framed as a linear sum assignment by assuming an injective mapping from the smaller graph to the larger one (Bonneel & Coeurjolly, 2019; Wang et al., 2019). A key advantage of this approach is that it can be efficiently solved with cubic worst-case time complexity using the Hungarian algorithm (Kuhn, 1955). However, Partial graph matching does not assume a total mapping, adding an extra layer of complexity. It requires selecting the subset of nodes to be matched while determining the mapping, making the problem more challenging. Consequently, traditional linear assignment methods like linear sum or sum-k assignment cannot be directly applied.
@@
Despite its challenging nature, recent studies have attempted to address the partial graph matching problem using deep learning techniques (Jiang et al., 2022; Wang et al., 2023a). One approach (Jiang et al., 2022) frames the problem as an Integer Linear Programming (ILP) task with dummy nodes. However, ILPs rely on branch and bound algorithms, which lack polynomial worst-case time complexity, and the use of dummy nodes can hinder node representation learning (Wang et al., 2023a). Another approach (Wang et al., 2023a) estimates the number of matchings (k) between two graphs, solving it as a $k$-assignment problem (Burkard et al., 2012) using GreedyTopK algorithm (Wang et al., 2023a). However, determining the optimal k is difficult and requires separate neural modules, leading to multiple training stages. These limitations highlight the need for an efficient optimization objective that captures the inherent nature of partial graph matching.
@@
In optimal partial transport formulations, divergence functions are typically used to penalize untransported mass, which in turn determines the amount of mass that will be transported. We theoretically show that using weighted total variation as the divergence function allows for an optimal partial transport objective that guarantees the existence of a partial assignment while incorporating node matching bias. Based on these theoretical properties, we define a new optimization problem for partial graph matching. We further demonstrate that this optimization is connected to the linear sum assignment problem, enabling the use of the Hungarian algorithm to exactly solve the partial graph matching problem.
@@
  • We define a new optimization problem for partial graph matching, inspired by optimal partial transport. By incorporating a cost matrix with matching biases, this formulation provides a robust optimization objective that carefully balances the selection of matched and unmatched nodes.
  • To solve the proposed optimization problem more efficiently, we further explore the underlying structure of its solution spaces for partial graph matching. This reveals a notable connection between partial graph matching and the linear sum assignment problem, allowing us to solve the partial graph matching problem exactly using the Hungarian algorithm.
@@ 2.1 Deep graph matching
Several graph matching methods leverage neural networks to learn matching-aware node embeddings (Jiang et al., 2022; Rolínek et al., 2020; Wang et al., 2020a; Yu et al., 2020). These methods integrate cross-graph node affinity with feature and structural data to learn node embeddings. The learned embeddings are then used to derive cross-graph node-to-node affinities. Then, the soft correspondences between nodes are typically obtained by applying Sinkhorn normalization on the cost-graph node-to-node affinity matrix. Once the soft correspondence matrix is obtained, it is projected into the space of permutation like binary matrices to achieve one-to-one node correspondences using algorithms such as the Hungarian algorithm (Kuhn, 1955) and Stable Matching algorithm (Ratnayaka et al., 2023), assuming a total injective mapping from the smaller graph to the larger graph.
So far, only a few deep learning approaches have addressed partial graph matching. (Wang et al., 2023b) proposes a method assuming a known number of node pairs ($k$), which limits its applicability.  (Wang et al., 2023a) employs separate neural networks to estimate match counts, but this increase complexity and error propagation. Another study  (Jiang et al., 2022) introduces dummy nodes to bypass explicit match number estimation, framing the problem as an Integer Linear Programming (ILP) task. However, ILP with branch and bound suffers from high time complexity, and the use of dummy nodes can distort node representations by implying higher similarity with unmatched nodes. The end-to-end learning mechanisms of these state-of-the-art partial graph matching techniques are guided by loss functions tailored to their specific solutions. For instance, in (Jiang et al., 2022), the ILP solution selects elements from the doubly stochastic affinity matrix for loss calculation. In  (Wang et al., 2023a), the number of matchings (k) between two graphs is first estimated, and the loss is then computed based on the probability of a node pair being among the top k assignments.
@@ 7 Partial Graph matching Architecture
$L_{bias}=\frac{1}{m}\sum_{i=1}^{m}\left[(M\mathbf{1}_{n})_{i}-\alpha_{i}\right]^{2}+\frac{1}{n}\sum_{j=1}^{n}\left[(M^{\intercal}\mathbf{1}_{m})_{j}-\beta_{j}\right]^{2}.$
Here, $\alpha_{i}$ close to 1 indicates a higher chance of node $i$ being matched, while a value close to 0 indicates a lower chance. Similarly, $\beta_{j}$ close to 1 indicates a higher chance of node $j$ being matched. The goal of $L_{bias}$ is to adjust the matching bias so that it approaches 0 for nodes that should remain unmatched according to the ground truth, and 1 for nodes that should be matched. Specifically, for any $i\in V_{S}$, $(M\mathbf{1}_{n})_{i}=0$ means node $i$ should remain unmatched, and for any $j\in V_{T}$, $(M^{\intercal}\mathbf{1}_{m})_{j}=0$ means node $j$ should remain unmatched.
The partial matching loss is defined as $L=L_{cost}+\lambda L_{bias}$, where $\lambda\in(0,1]$ is the regularization parameter.
@@ 8.1 Experimental Setup
In this task, we focus on image keypoint matching, which aims to find corresponding annotated keypoints between two given images. We use three image keypoint matching datasets that inherently contain outliers: 1) the Pascal VOC Keypoint with Berkeley annotations (Everingham et al., 2010), which includes keypoint-annotated images from 20 classes; 2) SPair-71k (Min et al., 2019), featuring 70,958 high-quality image pairs from Pascal VOC 2012 and Pascal 3D+, covering 18 classes; and 3) IMC PT SparseGM (Wang et al., 2023a), a recently proposed dataset specifically for partial graph matching.
Experimental Setting.  We followed the same experimental setting and evaluation mechanism as described in  (Wang et al., 2023a) to conduct experiments on the image keypoint matching datasets. Within this experimental setup, a pre-trained VGG16 model (Simonyan & Zisserman, 2014) is used to extract visual features of annotated image keypoints. Graphs are created using the extracted keypoint features following the same protocol as in (Wang et al., 2023a). We use the neural network proposed in GCAN (Jiang et al., 2022) as $NN_{\theta}$ to learn node embeddings and compute the cross-graph node-to-node affinity matrix.
Consistent with (Wang et al., 2023a), we report the matching F1-score as the evaluation metric. When reporting the results, we run 5 random starts and report the 95% confidence interval as the error bars, similarly to  (Wang et al., 2023a). We considered the following baselines: NGM-v2  (Wang et al., 2020b), GCAN (Jiang et al., 2022), AFAT-U (Jiang et al., 2022), and AFAT-I (Wang et al., 2023a). Our evaluation is based on the implementation and results reported in (Wang et al., 2023a) [^1] https://github.com/Thinklab-SJTU/ThinkMatch. In  Table 2, and  Table 2(left), GM Network denotes the neural network architecture that each of the given baselines used to learn node embeddings and to obtain the cross-graph node-to-node affinity matrix ($A$). PMH indicates the technique that is used to distinguish matching vs non-matching nodes and to obtain the final partial matching.
@@
GCAN OPGM 51.9 43.8 66.6 28.9 60.9 60.6 37.8 67.8 37.7 64.3 58.9 47.6 47.8 43.1 77.3 49.5 82.1 90.9 56.5 $\pm$ 0.4
GCAN OPGM-rs 52.4 43.7 66.5 30.6 61.4 60.2 38.5 68.2 39.0 65.9 58.9 50.6 48.0 44.8 78.6 49.1 82.3 92.1 57.3$\pm$ 0.4
@@
Dataset Name IMCPT 50 IMCPT 100 GM Network PMH reichstag sacre st peters mean reichstag sacre st peters mean NGM-v2 dummy 88.5 56.1 63.0 69.2±0.5 80.0 57.0 71.3 69.5±0.3 NGM-v2 AFAT-U 90.5 58.7 66.9 72.0±0.3 81.7 57.0 72.2 70.3±0.2 NGM-v2 AFAT-I 92.3 58.7 66.7 72.8±0.4 82.0 57.0 71.4 70.1±0.3 GCAN ILP 87.2 55.1 63.0 68.4±0.5 80.4 55.7 72.8 69.6±0.4 GCAN AFAT-U 86.9 59.4 67.1 71.1±0.4 82.6 58.2 73.8 71.5±0.2 GCAN AFAT-I 91.0 60.3 67.3 72.9±0.6 82.7 57.8 72.4 70.9±0.4 GCAN OPGM 91.2 57.0 68.1 71.9±0.2 81.7 59.1 76.1 72.3±0.4 GCAN OPGM-rs 92.0 58.6 67.8 72.8±0.1 82.1 59.5 75.9 72.5±0.3
OPGM-rsAFAT-IAFAT-UGCAN00100100100100200200200200300300300300400400400400500500500500600600600600117.5117.5117.5117.5367.1367.1367.1367.1371.8371.8371.8371.8561.1561.1561.1561.1(ms)Inference time
@@ 8.2 Results and Discussion
Table 2 (left) shows the results for the IMCPT 50 and IMCPT 100 datasets. Our models outperform the baselines in terms of mean F1-score on the IMCPT 100 dataset and in 2 out of 3 classes. For the IMCPT 50 dataset, our method performs comparably to the best baselines. Notably, the IMCPT 100 dataset, which contains the highest number of nodes among visual graph matching datasets (Wang et al., 2023a), further highlights the efficacy of our approach for partial graph matching. We discuss the evaluations on the Pascal VOC Keypoint dataset in the appendix, where we observe that poor and ambiguous annotations in data can adversely affect the performances of our proposed models.
@@
We conducted experiments to investigate the impact of the unbalancedness parameter, $\rho$, and the regularization parameter $\lambda$ on the partial graph matching performance. The experiments were performed on two datasets: Spair-71K, and IMCPT 100.
We first evaluated the performance of the OPGM-rs model with varying values of $\rho$, while keeping other hyperparameters fixed. Fig. 2 (middle) shows the mean F1-score for varying $\rho$, indicating its critical role in performance. Smaller $\rho$ values may reduce performance by restricting valid matches, while larger values of $\rho$ can also decrease performance as it may allow incorrect matches. We also examined the effect of the regularization parameter $\lambda$ in the partial matching loss. Starting from $0.01$, we evaluated the OPGM-rs model with varying $\lambda\in\{0.2,0.4,0.6,0.8,1\}$, while keeping other hyperparameters fixed. As shown in Fig. 2 (right), while there is slight variation in the mean F1-score as $\lambda$ changes, the impact is not significant.
@@
We evaluated the average runtime for processing a pair of graphs (i.e., keypoint-annotated images) during the inference phase for OPGM-rs, GCAN, AFAT-I, and AFAT-U using the IMCPT 100 dataset, the largest image keypoint matching dataset (Wang et al., 2023a). Matching was performed on 3000 keypoint-annotated pairs, and the average time per sample was calculated. As shown in Table 2 (right), OPGM-rs demonstrates significantly lower inference time compared to the other models. It is important to note that all models use the same neural network architecture from  (Jiang et al., 2022) (corresponding to $\textsc{nn}_{\theta}$) to compute the cross-graph node-to-node affinity matrix $A$ and all implement the Sinkhorn algorithm to derive the doubly stochastic affinity matrix $S$. The main difference between these models lies in the technique used to obtain the partial matching between the two node sets.

### v4

# Learning Partial Graph Matching via Optimal Partial Transport

@@ Abstract
Partial graph matching extends traditional graph matching by allowing some nodes to remain unmatched, enabling applications in more complex scenarios. However, this flexibility introduces additional complexity, as both the subset of nodes to match and the optimal mapping must be determined. While recent studies have explored deep learning techniques for partial graph matching, a significant limitation remains: the absence of an optimization objective that fully captures the problem’s intrinsic nature while enabling efficient solutions. In this paper, we propose a novel optimization framework for partial graph matching, inspired by optimal partial transport. Our approach formulates an objective that enables partial assignments while incorporating matching biases, using weighted total variation as the divergence function to guarantee optimal partial assignments. Our method can achieve efficient, exact solutions within cubic worst case time complexity. Our contributions are threefold: (i) we introduce a novel optimization objective that balances matched and unmatched nodes; (ii) we establish a connection between partial graph matching and linear sum assignment problem, enabling efficient solutions; (iii) we propose a deep graph matching architecture with a novel partial matching loss, providing an end-to-end solution. The empirical evaluations on standard graph matching benchmarks demonstrate the efficacy of the proposed approach.
@@ 1 Introduction
Graph matching is a fundamental problem in network analysis, aiming to establish one-to-one correspondences between nodes in two graphs based on a defined objective. It has broad applications across various fields, including computer vision (Sun et al., 2020), bioinformatics (Zaslavskiy et al., 2009), and social network analysis (Zhang et al., 2019), where it is used to solve complex real-world problems. Traditional graph matching assumes a bijective mapping between the nodes of two graphs or a total injective mapping from the smaller graph to the larger one. However, these assumptions often restrict its applicability in more complex, real-world scenarios. Partial graph matching, a generalized version of the graph matching problem, addresses these limitations by allowing some nodes in both graphs to remain unmatched (Wang et al., 2023; Jiang et al., 2022b). It seeks an optimal partial assignment, an injective partial function between node sets, thereby expanding the practical utility of graph matching. This approach is particularly useful in applications where not all nodes have meaningful counterparts. For example, in image keypoint matching, not all keypoints in two given images have correspondences (Jiang et al., 2022b), and in biological networks, certain proteins may lack direct counterparts in other species (Zaslavskiy, 2010).
@@
Solving partial graph matching is inherently more complex than traditional graph matching due to the additional challenge of determining both the subset of nodes to match and the optimal mapping itself. Graph matching is generally framed as a combinatorial optimization problem, where the choice of optimization objective is crucial. Early works formulated it as a Quadratic Assignment Problem (QAP), which is NP-hard (Koopmans & Beckmann, 1957; Lawler, 1963). Recent approaches  (Wang et al., 2021; 2020a; Rolínek et al., 2020; Fey et al., 2020) use neural networks to learn node representations and derive a cross-graph node-to-node cost matrix. This allows graph matching to be framed as a linear sum assignment problem (Burkard et al., 2012; Yu et al., 2020), assuming a bijective mapping between two node sets of equal size. Even if two graphs have an unequal number of nodes, the problem can still be framed as a linear sum assignment by assuming an injective mapping from the smaller graph to the larger one (Bonneel & Coeurjolly, 2019; Wang et al., 2019b). A key advantage of this approach is that it can be efficiently solved with cubic worst-case time complexity using the Hungarian algorithm (Kuhn, 1955). However, Partial graph matching does not assume a total mapping, adding an extra layer of complexity. It requires selecting the subset of nodes to be matched while determining the mapping, making the problem more challenging. Consequently, traditional linear assignment methods like linear sum or k assignment cannot be directly applied.
@@
Despite its challenging nature, recent studies have attempted to address the partial graph matching problem using deep learning techniques (Jiang et al., 2022b; Wang et al., 2023). One approach (Jiang et al., 2022b) frames the problem as an Integer Linear Programming (ILP) task with dummy nodes. However, ILPs rely on branch and bound algorithms, which lack polynomial worst-case time complexity, and the use of dummy nodes can hinder node representation learning (Wang et al., 2023). Another approach (Wang et al., 2023) estimates the number of matchings (k) between two graphs, solving it as a $k$-assignment problem (Burkard et al., 2012) using GreedyTopK algorithm (Wang et al., 2023). However, determining the optimal k is difficult and requires separate neural modules, leading to multiple training stages. These limitations highlight the need for an efficient optimization objective that captures the inherent nature of partial graph matching.
@@
In optimal partial transport formulations, divergence functions are typically used to penalize untransported mass, which in turn determines the amount of mass that will be transported. We theoretically show that using weighted total variation as the divergence function allows for an optimal partial transport objective that guarantees the existence of a partial assignment while incorporating node matching bias. Based on these theoretical properties, we define a new optimization problem for partial graph matching. We further demonstrate that the partial matching can solved by embedding it in a linear assignment problem. This enables the use of the Hungarian algorithm to exactly solve the partial graph matching problem.
@@
  • We define a new optimization problem for partial graph matching, inspired by optimal partial transport. By incorporating a cost matrix as well as matching biases, this formulation provides a robust optimization objective that carefully balances the selection of matched and unmatched nodes.
  • To solve the proposed optimization problem efficiently, we explore the underlying structure of its solution spaces. This reveals a notable embedding of the partial graph matching problem in a linear sum assignment problem. Furthermore, the solutions of the latter assignment problem can be mapped efficiently to solutions of the partial matching problem, which are themselves optimal. This allows us to solve the partial graph matching problem exactly using the Hungarian algorithm.
@@ 2.1 Deep graph matching
Several graph matching methods leverage neural networks to learn matching-aware node embeddings (Jiang et al., 2022b; Rolínek et al., 2020; Wang et al., 2020a; Yu et al., 2020; Gao et al., 2021). These methods integrate cross-graph node affinity with feature and structural data to learn node embeddings. The learned embeddings are then used to derive cross-graph node-to-node affinities. Then, the soft correspondences between nodes are typically obtained by applying Sinkhorn normalization on the cost-graph node-to-node affinity matrix. Once the soft correspondence matrix is obtained, it is projected into the space of permutation like binary matrices to achieve one-to-one node correspondences using algorithms such as the Hungarian algorithm (Kuhn, 1955) and Stable Matching algorithm (Ratnayaka et al., 2023), assuming a total injective mapping from the smaller graph to the larger graph.
Optimal transport techniques have also been discussed to solve the graph matching problem  (Xu et al., 2019), where each node in the source graph is matched to a node in the target graph. Moreover, optimal partial transport techniques have been applied to subgraph matching  (Pan et al., 2024), where a preset fraction of mass (corresponding to a fixed number of nodes) from a smaller graph is matched to nodes in a larger graph. This approach constrains the matching process by specifying in advance how many nodes from the smaller graph must be matched. However, these approaches are not directly applicable to partial graph matching, which requires identifying corresponding nodes between two graphs without any prior assumptions about the number of nodes to be matched. Moreover,spectral methods have been successfully applied to both graph matching  (Wang et al., 2019a) and also to compactly encode maps between graphs and subgraphs  (Pegoraro et al., 2022). While these problems share similarities in attempting to find correspondences at node or graph level, they differ fundamentally from partial graph matching in their matching objectives and constraints.
So far, only a few deep learning approaches have addressed partial graph matching. Jiang et al. (2022b) introduces dummy nodes to bypass explicit match number estimation, framing the problem as an Integer Linear Programming (ILP) task. However, ILP with branch and bound suffers from high time complexity, and the use of dummy nodes can distort node representations by implying higher similarity with unmatched nodes. Wang et al. (2023) attempts to solve partial graph matching as a k-assignment problem. However, as the match count (k) between two graphs is not known in a partial graph matching problem, their approach to handle partial matchings include two steps, first estimating the match count (k) by using a separate neural module and solving an entropic optimal transport problem and then using the estimated k value to solve partial graph matching as a k-assignment problem. This two step approach increase computational complexity and error propagation. Another line of work  (Nurlanov et al., 2023; Jiang et al., 2022a) to perform partial graph matching have explored universe graph representation learning, where a "universe" graph is constructed for each object class in keypoint-based image analysis. However, these methods require prior knowledge specific to each class, such as the exact number of distinct keypoints (nodes) in a class, making them difficult to generalize to graphs or images from previously unseen classes. In contrast, our work addresses a more general form of partial graph matching problem that operates without any class-specific assumptions, allowing it to work with arbitrary graph structures.
@@ 7 Partial Graph matching Architecture
$L_{bias}=\frac{1}{m}\sum_{i=1}^{m}(M\mathbf{1}_{n})_{i}\left[1-\alpha_{i}\right]^{2}+\frac{1}{n}\sum_{j=1}^{n}(M^{\intercal}\mathbf{1}_{m})_{j}\left[1-\beta_{j}\right]^{2}.$
The goal of $L_{bias}$ is to increase the matching bias of a node that should be matched. The partial matching loss is defined as $L=L_{cost}+\lambda L_{bias}$, where $\lambda\in(0,1]$ is the regularization parameter.
@@ 8.1 Experimental Setup
In this task, we focus on image keypoint matching, which aims to find corresponding annotated keypoints between two given images. We use three image keypoint matching datasets that inherently contain outliers: 1) the Pascal VOC Keypoint with Berkeley annotations (Everingham et al., 2010), which includes keypoint-annotated images from 20 classes; 2) SPair-71k (Min et al., 2019), featuring 70,958 high-quality image pairs from Pascal VOC 2012 and Pascal 3D+, covering 18 classes; and 3) IMC PT SparseGM (Wang et al., 2023), a recently proposed dataset specifically for partial graph matching.
Experimental Setting.  We followed the same experimental setting and evaluation mechanism as described in  (Wang et al., 2023) to conduct experiments on the image keypoint matching datasets. Within this experimental setup, a pre-trained VGG16 model (Simonyan & Zisserman, 2014) is used to extract visual features of annotated image keypoints. Graphs are created using the extracted keypoint features following the same protocol as in (Wang et al., 2023). We use the neural network proposed in GCAN (Jiang et al., 2022b) as $NN_{\theta}$ to learn node embeddings and compute the cross-graph node-to-node affinity matrix.
Consistent with (Wang et al., 2023), we report the matching F1-score as the evaluation metric. When reporting the results, we run 5 random starts and report the 95% confidence interval as the error bars, similarly to  (Wang et al., 2023). We considered the following baselines: NGM-v2  (Wang et al., 2020b), GCAN (Jiang et al., 2022b), AFAT-U (Jiang et al., 2022b), and AFAT-I (Wang et al., 2023). Our evaluation is based on the implementation and results reported in (Wang et al., 2023) [^1] https://github.com/Thinklab-SJTU/ThinkMatch. In  Table 2, and  Table 2(left), GM Network denotes the neural network architecture that each of the given baselines used to learn node embeddings and to obtain the cross-graph node-to-node affinity matrix ($A$). PMH indicates the technique that is used to distinguish matching vs non-matching nodes and to obtain the final partial matching.
@@
GCAN OPGM 51.9 43.8 66.6 28.9 60.9 60.6 37.8 67.8 37.7 64.3 58.9 47.6 47.8 43.1 77.3 49.5 82.1 90.9 56.5 ± 0.4
GCAN OPGM-rs 52.5 43.2 67.9 32.3 61.2 61.2 39.9 68.9 38.2 67.1 60.0 50.1 47.4 46.3 78.8 50.2 82.9 92.7 57.8 ± 0.2
@@
Dataset Name IMCPT 50 IMCPT 100 GM Network PMH reichstag sacre st peters mean reichstag sacre st peters mean NGM-v2 dummy 88.5 56.1 63.0 69.2±0.5 80.0 57.0 71.3 69.5±0.3 NGM-v2 AFAT-U 90.5 58.7 66.9 72.0±0.3 81.7 57.0 72.2 70.3±0.2 NGM-v2 AFAT-I 92.3 58.7 66.7 72.8±0.4 82.0 57.0 71.4 70.1±0.3 GCAN ILP 87.2 55.1 63.0 68.4±0.5 80.4 55.7 72.8 69.6±0.4 GCAN AFAT-U 86.9 59.4 67.1 71.1±0.4 82.6 58.2 73.8 71.5±0.2 GCAN AFAT-I 91.0 60.3 67.3 72.9±0.6 82.7 57.8 72.4 70.9±0.4 GCAN OPGM 91.2 57.0 68.1 71.9±0.2 81.7 59.1 76.1 72.3±0.4 GCAN OPGM-rs 91.7 59.2 67.2 72.7±0.1 82.6 59.6 75.9 72.7±0.2
OPGM-rsAFAT-IAFAT-UGCAN00100100100100200200200200300300300300400400400400500500500500600600600600117.8117.8117.8117.8367.1367.1367.1367.1371.8371.8371.8371.8561.1561.1561.1561.1(ms)Inference time
@@ 8.2 Results and Discussion
Table 2 (left) shows the results for the IMCPT 50 and IMCPT 100 datasets. Our models outperform the baselines in terms of mean F1-score on the IMCPT 100 dataset and in 2 out of 3 classes. For the IMCPT 50 dataset, our method performs comparably to the best baselines. Notably, the IMCPT 100 dataset, which contains the highest number of nodes among visual graph matching datasets (Wang et al., 2023), further highlights the efficacy of our approach for partial graph matching. We discuss the evaluations on the Pascal VOC Keypoint dataset in the appendix, where we observe that poor and ambiguous annotations in data can adversely affect the performances of our proposed models.
@@
We evaluated the impact of the unbalancedness parameter $\rho$ and the regularization parameter $\lambda$ on partial graph matching performance using Spair-71K and IMCPT 100 datasets. For $\rho$, we varied its values while keeping other hyperparameters fixed. As shown in Fig. 2 (middle), mean F1-scores highlight $\rho$’s critical role: smaller values restrict valid matches, while larger values may allow incorrect matches, both reducing performance. For $\lambda$, starting from $0.01$, we tested $\lambda\in\{0.2,0.4,0.6,0.8,1\}$ with fixed hyperparameters. As shown in Fig. 2 (right), the mean F1-scores show slight variation, indicating that the impact of $\lambda$ is not significant.
@@
We evaluated the average runtime for processing a pair of graphs (i.e., keypoint-annotated images) during the inference phase for OPGM-rs, GCAN, AFAT-I, and AFAT-U using the IMCPT 100 dataset, the largest image keypoint matching dataset (Wang et al., 2023). Matching was performed on 3,000 keypoint-annotated pairs. As shown in Table 2 (right), OPGM-rs demonstrates significantly lower inference time compared to the other models. It is important to note that all models use the same neural network architecture from  (Jiang et al., 2022b) (corresponding to $\textsc{nn}_{\theta}$) to compute the cross-graph node-to-node affinity matrix $A$ and the Sinkhorn algorithm to derive the doubly stochastic affinity matrix $S$, differing only in their partial matching techniques.